### v1

# Centrality in Collaboration: A Novel Algorithm for Social Partitioning Gradients in Community Detection for Multiple Oncology Clinical Trial Enrollments

Abstract
Patients at a comprehensive cancer center who do not achieve cure or remission following standard treatments often become candidates for clinical trials. Patients who participate in a clinical trial may be suitable for other studies. A key factor influencing patient enrollment in subsequent clinical trials is the structured collaboration between oncologists and most responsible physicians. Possible identification of these collaboration networks can be achieved through the analysis of patient movements between clinical trial intervention types with social network analysis and community detection algorithms. In the detection of oncologist working groups, the present study evaluates three community detection algorithms: Girvan-Newman, Louvain and an algorithm developed by the author. Girvan-Newman identifies each intervention as their own community, while Louvain groups interventions in a manner that is difficult to interpret. In contrast, the author’s algorithm groups interventions in a way that is both intuitive and informative, with a gradient evident in social partitioning that is particularly useful for epidemiological research. This lays the groundwork for future subgroup analysis of clustered interventions.
1 Introduction
When cancer patients complete standard treatments at Princess Margaret Cancer Centre and have not responded with being cured or in remission, they become candidates for clinical trials. These clinical trials are regulated studies registered by Health Canada [^1] For more information, see https://www.canada.ca/en/health-canada/services/clinical-trials.html as opposed to quality assurance studies [^2] Quality assurance studies in the context of medical studies are studies which look at drugs which are already approved for use, but the goals are focused on other aspects of care such as drug delivery or quality of care.. Patients who qualify may have been screen failures for other trials, have experienced progressive disease, or are receiving maintenance therapy and have been referred to a clinical trial by their oncologist or most responsible physician. Ground truth shows that collaboration networks between oncologists is a primary factor for further engagement in subsequent clinical trials by patients after completion of the given clinical trial that they are enrolled in. A possible approach to understanding the structure of these collaboration networks is through use of social network analysis (SNA) and community detection algorithms.
Social network analysis examines individual entities and their relationships among them. The data is represented as a “graph” where individual entities are referred to as “nodes” and their relationships between them as “edges”, which may be directional if specified (see Figure 1). A primary area of study in SNA is the analysis of interconnectivity of nodes, called ”communities” and identification of clusters through the use of algorithms called ”community detection algorithms”. Rostami et al1 (2023) note that there is no specific model which describes exactly what a ”community” is. Generally, community detection algorithms employ specific optimization strategies to partition a large-scale complex network into a set of disjoint and compact subgroups, often (but not always) without prior knowledge regarding the number of subgroups and their sizes. Rostami et al further note that it is commonly acknowledged that there is no unique community detection algorithm that can accommodate all kinds of graphs, because of the inherent variability in network structures and their respective objective(s).
Figure 1: Two simple graphs with directed and undirected edges. Direction is noted by arrowheads at the end of the edges.
Application of community detection algorithms with oncology clinical trial data has been preformed in the past. Georgiev et al2 (2011) applied the Girvan-Newman3 (2002) algorithm and noted a lack of cohesion among researchers who studied treatments for multiple myeloma. Haq and Wang4 (2016) applied the Louvain algorithm (by Blondel et al5 (2008)) to identify communities of cancer patients with significantly different survival curves. The present study applies SNA, and compares multiple community detection algorithms to identify collaboration networks between oncologists through the interventions studied in clinical trials via enrollment data of patients in multiple, nonconcurrent clinical trials at Princess Margaret Cancer Centre (PM) in Toronto, Canada. Inspired by work from Gorgiev et al (2011), Haq and Wang (2016), Ostovari and Yu6 (2019) and Bissoyi and Patra7 (2020) this research considers the Girvan-Newman and Louvain algorithms and compares them to an author-developed algorithm, referred to as “Smith-Pittman” [^3] Named after the author and his co-supervisor, Tyler Pittman., to identify collaboration networks between clinical trials classified by intervention.
2 Materials and Methods
2.1 The Data
The data is a result of multiple data source integration, which was undertaken from the PM Cancer Registry and Clinical Research Record. The data is anonymized, and spans patient enrollments in oncology clinical trials between January 1, 2016 and December 31, 2018. In this time period, there were 2970 patients enrolled in 515 clinical trials involving 41 principal investigators. For the identification of collaboration networks between oncologists, the analytic sample only consists of patients who were enrolled in more than one clinical trial within the time period studied. The resulting analytic sample consists of 389 patients enrolled in 288 clinical trials. Among these clinical trials, some interventions can be classified into broader categories of targeted therapies, or immunotherapy. This has been identified in the data with “T:” and “I:” prefixes respectively. The clinical trials were classified by intervention type, presenting as 16 distinct intervention types among 470 patient enrollments. With this classification, the patient referral graph is constructed (see Figure 2).
The analysis is preformed with the R programming language, and makes use of an extensive array of libraries and dependencies. The primary libraries that were utilized include igraph, tidyverse, and tidygraph. For the complete script, please refer to the Appendix - Program Syntax.
Figure 2: Patient movement between clinical trials classified by intervention type at PM. Nodes indicate the treatment type, and labeled edges indicate the movement (subsequent enrollment) of patients between clinical trials in a given intervention of the same type (self loop), or differing. Among the clinical trials, some interventions can be classified into broader categories consisting of targeted therapies or immunotherapy. This has been identified in the data with “T:” and “I:” prefixes respectively.
2.2 Methods
The goal of applying community detection algorithms with this data is to identify oncologist working groups among treatment interventions, based on the movement (incoming and outgoing referrals) of patients between the intervention types. These movements in the network are understood through measures that are considered by the community detection algorithms’ optimization strategies. While the Girvan-Newman, Louvain and Smith-Pittman algorithms differ in their approaches to the identification of collaboration networks, their identification strategies are based on the maximization of modularity, $Q$- a measure that scores the degree of segregation within a network through tightly connected communities or clusters (See Newman8 (2006)).
The mathematical representation of modularity is defined in the igraph R package9 (2006) as:
$Q=\frac{1}{m}\sum_{i,j}\left(A_{ij}-\frac{k_{i}^{\text{out}}k_{j}^{\text{in}}}{m}\right)\delta\left(c_{i},c_{j}\right)$
Where $m$ is the number of edges (patient movements), $A_{ij}$ is the number of connections shared by nodes $i$ and $j$ (movements between interventions $i$ and $j$), $k_{i}^{\text{out}}$ and $k_{j}^{\text{in}}$ are the number of edges coming out from node $i$ and going into node $j$ (patient movements from intervention $i$ and $j$) and $\delta\left(c_{i},c_{j}\right)$ is an indicator variable identifying if nodes $i$ and $j$ are connected - either directly or through another node (if there is a patient movement between interventions $i$ and $j$ either directly or through some other intervention). For directed graphs, $k_{i}^{\text{out}}$ and $k_{j}^{\text{in}}$ are simply the number of connected edges possessed by nodes $i$ and $j$, respectfully. For a more comprehensive overview modularity and other measures in social network analysis, see Newman (2006), Wasserman and Faust10 (1994) and Latora et al11 (2017).
2.2.1 Girvan-Newman
The Girvan-Newman algorithm is based on the evaluation of edges in a social network through edge-betweenness centrality. Edge-betweenness centrality is defined by Girvan and Newman (2002) as the number of shortest paths that go through an edge in a graph, divided by the total number of shortest paths between nodes $i$ and $j$. Each edge in a graph has its own edge-betweenness centrality value. The igraph (2006) documentation defines edge-betweenness centrality for an edge $e$ in a social network in mathematical terms as:
$\sum_{i\neq j}g_{iej}/g_{ij}$
Where $g_{ij}$ is the number of shortest paths between nodes $i$ and $j$ (patient movements between interventions $i$ and $j$, either directly or through some other intervention(s)), and $g_{iej}$ is the number of shortest paths which pass through edge $e$. Figure 3 provides an illustration of a simple network, showing the edge with the highest edge-betweenness centrality highlighted in red.
Figure 3: A simple network demonstrating an edge with a high edge-betweenness centrality, highlighted in red. The network consists of two densely connected clusters, with the red edge serving as the sole connection between them. This edge is crucial for communication between the two clusters, as most of the shortest paths that connect nodes from opposite clusters pass through it.
Edge betweenness can be calculated for directed and undirected edges. As a result, the Girvan-Newman algorithm can be applied to directed or undirected graphs without any transformations. The steps of the Girvan-Newman algorithm are as follows:
  1. 1.
    Calculate edge-betweenness centrality for all edges in the network.
  2. 2.
    Find the edge with the highest edge-betweenness centrality, and remove it from the network.
  3. 3.
    Recalculate edge-betweenness centrality for all remaining edges.
  4. 4.
    Repeat from step 2.
Girvan-Newman can be utilized when the community structure is known, and will classify nodes into a predetermined number of communities based on the hierarchy produced by the algorithm (see Girvan and Newman 2002). When the community structure is not known, modularity is evaluated after each iteration of the algorithm. The grouping of nodes into distinct communities is selected via modularity maximization.
2.2.2 Louvain
The Louvain algorithm (by Blondel et al 2008) operates in two distinct phases. (i) In the first phase, each node in the network is considered as their own community, resulting in the initial partition with as many communities as there are nodes. The algorithm then assesses the potential modularity gain for each node $i$ if it were to leave its current community and join the community of node $j$. After evaluating the potential modularity gain across all communities, node $i$ is reassigned to the community of node $j$, where the modularity increase is maximized. The process is iteratively and sequentially applied for all nodes, until no further improvement can be achieved. This first phase stops when a local maximum of modularity is reached, meaning that no subsequent node move can enhance modularity. (ii) The second phase involves constructing a new network as represented by the communities identified in the first phase. Links between nodes of the same community are viewed as “self-loops” for the community in the new network. Once this second phase is complete, the first phase of the algorithm can be reapplied. The combination of these two phases is referred to as a “pass”. The algorithm terminates when there is no other local maxima in modularity to be achieved in subsequent passes.
A key limitation of the Louvain algorithm is that it is generally programmed to work only with undirected graphs [^4] Work on extending the Louvain algorithm to accommodate directed graphs has been an outstanding issue in the igraph developer community since 2015 (See: https://github.com/igraph/igraph/issues/890). However, Dugué and Perez12 (2022) have done some work on this.. In order to apply the Louvain algorithm to a directed graph, it must first be converted to an undirected graph. Figure 4 is a reproduction of Blodel et al’s (2008) illustration of the algorithm.
Figure 4: Reproduced illustration of the Louvain algorithm (originally designed by Blondel et al (2008)).
2.2.3 Smith-Pittman
The ”Smith-Pittman” algorithm is a modification of the Girvan-Newman algorithm, where degree centrality is considered. Degree centrality of a node is simply defined as the number of connections a node has within a given network (see figure 5). The algorithm proceeds through the following steps:
  1. 1.
    Calculate the degree centrality for each node, and the edge-betweenness centrality of all edges in the network.
  2. 2.
    Identify the subgraph associated with the node that has the highest degree centrality.
  3. 3.
    Remove the edge possessing the highest calculated edge-betweenness centrality.
  4. 4.
    Recalculate the degree centrality for all nodes, and the edge-betweenness centrality for the remaining edges in the network.
  5. 5.
    Repeat from step 2.
Figure 6 provides a visual representation of this algorithm. Like Girvan-Newman, the Smith-Pittman algorithm can be applied to both directed and undirected graphs. Conceptually, the algorithm can be specified to terminate once a predetermined number of communities have been identified. However, its primary design is for use in an unsupervised setting, where clusters are identified through the maximization of modularity as evaluated after each iteration of the algorithm.
Figure 5: A simple network highlighting node degree. The center node (colored red) possesses the highest number of connections and as a result possesses the highest degree and degree centrality index.
Figure 6: Illustration of the Smith-Pittman algorithm. Highlighted convex hulls denote the identification of distinct communities.
3 Results
Figures 7-9 show the communities identified by the algorithms, through convex hulls highlighting the grouped interventions. Tables 1-3 show the grouping of interventions into communities, and the breakdown by frequency of incoming and outgoing patient referrals for each treatment intervention studied. Figure 7 demonstrates that the Girvan-Newman algorithm identified each intervention as a separate community ($Q=$ 0.044). This result is particularly uninformative, as it is equivalent to not applying any community detection method to identify oncologist collaboration networks between the interventions. Figure 8 shows that the Louvain algorithm groups interventions into four distinct working groups, achieving the highest modularity score ($Q=$ 0.177). However, the underlying rationale and meaning behind these groupings remains unclear, beyond the objective to cluster interventions as to maximize modularity.
Figure 9 shows that the Smith-Pittman algorithm ($Q=$ 0.08) identified eight communities. Six of these communities consist of individual interventions - namely T: Small Molecule, I:MAbs [^5] Short for Monoclonal Antibodies. Checkpoint, I:Combined, I:MAbs Targeting, Combined Modality and Radioconjugate - while the remaining two communities encompass multiple interventions. The interpretation of the communities identified by the Smith-Pittman algorithm can be facilitated by the degree of connectivity among the interventions within these communities. Communities comprised of individual interventions either have the highest or a substantial number of patient referrals, whether incoming from or outgoing to other interventions, or they have the least. Figure 10 illustrates the distribution of interventions by patient referrals, ordered from smallest to largest, and highlights the thresholds beyond which single intervention communities are positioned. The interpretation of the communities identified by the Smith-Pittman algorithm suggests the existence of both highly connected, and less connected interventions, as well as broader groups corresponding to typical intervention types - there is a gradient that is evident in social partitioning. This interpretation offers an intuitive understanding related to the formation of collaboration networks being a function of intervention “popularity” - i.e. patient referrals outgoing and incoming to and from other interventions.
Figure 7: Detected communities via Girvan-Newman with modularity maximization. Sixteen distinct communities.
Intervention Referrals In Referrals Out Total
Chemotherapy 4 10 14
Combined Modality 0 1 1
Drug Repurposing 7 3 10
I: Adoptive Cell Transfer 10 3 13
I: Anti Cancer Vaccine 4 7 11
I: Combined 54 22 76
I: MAbs Checkpoint 92 147 239
I: MAbs Co-Stimulatory 31 22 53
I: MAbs Targeting 31 22 53
I: Oncolytic Virus 4 5 9
I: Other 25 6 31
Radioconjugate 1 0 1
T: Antibody-Drug Conjugate 18 10 28
T: Combined 9 8 17
T: Monoclonal Antibody 6 16 22
T: Small Molecule 174 188 362
Table 1: Girvan-Newman communities identified. Each intervention is their own community.
Figure 8: Detected communities via Louvain algorithm with modularity maximization. Four distinct communities.
Intervention Referrals In Referrals Out Total
Community: 1
Drug Repurposing 7 3 10
I: Other 25 6 31
T: Monoclonal Antibody 6 16 22
T: Small Molecule 174 188 362
Community: 2
Combined Modality 0 1 1
I: Adoptive Cell Transfer 10 3 13
I: Anti Cancer Vaccine 4 7 11
I: MAbs Checkpoint 92 147 239
Radioconjugate 1 0 1
T: Antibody-Drug Conjugate 18 10 28
Community: 3
I: Combined 54 22 76
I: MAbs Co-Stimulatory 31 22 53
I: Oncolytic Virus 4 5 9
Community: 4
Chemotherapy 4 10 14
I: MAbs Targeting 31 22 53
T: Combined 9 8 17
Table 2: Louvain communities identified and grouped interventions
Figure 9: Detected communities via Smith-Pittman algorithm with modularity maximization. Eight distinct communities.
Figure 10: Referral distribution among interventions. Interventions outside the boundaries (T: Small Molecule, I:MAbs Checkpoint, I: Combined, I:Mabs Targeting, Radioconjugate and Combined Modality) are each identified as individual communities, while interventions within them are identified as belonging to communities consisting of multiple interventions.
Intervention Referrals In Referrals Out Total
Community: 1
T: Small Molecule 174 188 362
Community: 2
I: MAbs Checkpoint 92 147 239
Community: 3
Chemotherapy 4 10 14
I: Adoptive Cell Transfer 10 3 13
I: Anti Cancer Vaccine 4 7 11
I: MAbs Co-Stimulatory 31 22 53
I: Oncolytic Virus 4 5 9
T: Antibody-Drug Conjugate 18 10 28
T: Combined 9 8 17
Community: 4
I: Combined 54 22 76
Community: 5
I: MAbs Targeting 31 22 53
Community: 6
Drug Repurposing 7 3 10
I: Other 25 6 31
T: Monoclonal Antibody 6 16 22
Community: 7
Combined Modality 0 1 1
Community: 8
Radioconjugate 1 0 1
Table 3: Smith-Pittman communities and identified and grouped interventions
4 Discussion
Where the Girvan-Newman algorithm failed to identify communities, the Louvain and Smith-Pittman algorithms succeeded. A possible explanation for this discrepancy lies in the nature of the data analyzed, which includes patient referrals to clinical trials that investigate the same intervention types as the clinical trials patients were previously enrolled in. In graph theory, such referrals are represented as “self loops” and introduce complexity in the network. The Girvan-Newman algorithm - whose original design was not for complex networks - struggles in such contexts, leading to its failure to group multiple interventions into communities based on modularity maximization.
The Louvain algorithm successfully detected communities. However, the resulting groups were difficult to interpret. This difficulty arises because the Louvain algorithm bases its community selection purely on modularity maximization, and does not consider the direction of patient movements the underlying structural or functional significance of particular interventions in the context of the network. The primary advantage of the Louvain algorithm is its efficiency in preforming community detection on large networks. It has been widely used in applications such as the Twitter Social Network (Pujol et al. 2009) which consisted of 2.4 million nodes and 38 million links, and mobile phone network data (Greene et al. 2010) with 4 million nodes, 100 million links. These networks are orders of magnitude larger than the patient referral network analyzed in this study, highlighting the scalability of the Louvain algorithm. However, utility of such a algorithm is limited in smaller, more specialized networks where the interpretability and justification of communities identified is important.
In contrast, the Smith-Pittman algorithm directly addresses connectivity of interventions studied in the clinical trials, by incorporating degree centrality and edge-betweenness centrality. This approach allows for the identification of communities with a more ordered structure, distinguishing between highly connected and minor interventions as they reflect the relational dynamics in the network. The results from the Smith-Pittman algorithm are promising, however, the results from this analysis alone is insufficient to establish generalizability of the algorithm. To fully assess its usefulness, a formal simulation study with further application of the Smith-Pittman algorithm in diverse settings is necessary. Additionally, the practical value of identified communities will become evident when they are applied as grouping variables in downstream analysis, such as outcome prediction or intervention effectiveness studies.
Further research should focus on subgroup analysis, and exploring extensions back to traditional statistical methods, such as regression and survival analysis. This research can further validate the utility of the identified communities, and the use of SNA and community detection algorithms in clinical research settings. The results of the Smith-Pittman algorithm lay the groundwork for these efforts, and potentially offer a robust tool for community detection in social and complex networks. Further work with the identified communities should involve assessment of the impact of community structure on patient outcomes, and identify if there are any structural inequities present in clinical trial enrollments. This line of research can lead to the identification of collaboration networks that improve patient care in clinical settings.